OpenStudy (rayman5766):

Which circles have an area of 16TT in2? Choose all answers that are correct. A. a circle with a radius of 2 in. B. a circle with a radius of 4 in. C. a circle with a diameter of 8 in. D. a circle with a diameter of 16 in

OpenStudy (rayman5766):

I think my answer is b and c

OpenStudy (igreen):

That is correct.

OpenStudy (igreen):

\(A = \pi r^2\) \(A = \pi (4^2)\) \(A = 16 \pi\) A diameter of 8 is the same as a radius of 4.

OpenStudy (rayman5766):

thanks iGreen I don't know what the other guy was doing

OpenStudy (igreen):

Lol